# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 2.1. G-Higgs bundles and Flat G-bundles
where $a$ is the co-action map. Here is an equivalent definition by Lie algebroids. Let ${\rm At}(V)$ be the Atiyah algebroid of $V$ (see [Bis10] or [CZ15, Appendix]). There is a natural projection ${\rm At}(V)\rightarrow T_{X}$ together with a short exact sequence
$0\rightarrow V(\mathfrak{g})\rightarrow{\rm At}(V)\rightarrow T_{X}\rightarrow 0$
A $G$-connection $\nabla:T_{X}\rightarrow{\rm At}(V)$ is a splitting of the exact sequence, and it is integrable if the splitting is a homomorphism of Lie algebroids.
Definition 2.2.
A flat $G$-bundle is a pair $(V,\nabla)$, where $V$ is a $G$-bundle and $\nabla$ is an integrable $G$-connection.
Example 2.3.
This example is discussed in [CZ15, Appendix]. Since it is very helpful to understand Higgs fields and connections locally, we also include it here. Let $V$ be a trivial $G$-bundle on $X$, and the Atiyah algebroid ${\rm At}(V)\cong V(\mathfrak{g})\oplus T_{X}$ is a direct sum. Then a connection $\nabla$ is equivalent to a morphism $T_{X}\rightarrow V(\mathfrak{g})$, and therefore, the sheaf of integrable connections on $X$ is isomorphic to a subsheaf of $V(\mathfrak{g})\otimes\Omega_{X}$ and denote it by $(V(\mathfrak{g})\otimes\Omega_{X})^{cl}$, where the upperscript is for the word “closed”. The corresponding element of a given connection in $(V(\mathfrak{g})\otimes\Omega_{X})^{cl}$ is usually regarded as the connection form.
Now let $V$ be an arbitrary $G$-bundle. We can find a covering $\mathcal{U}=\{U_{i}\}_{i\in I}$, where $I$ is the index set, such that $V|_{U_{i}}\cong G\times U_{i}$. Therefore, an integrable $G$-connection $\nabla$ on $V$ can be regarded as a section of $(V|_{U_{i}}(\mathfrak{g})\otimes\Omega_{U_{i}})^{cl}$ locally. This approach shows that given a Higgs field $\theta$ and an integrable $G$-connection $\nabla$, the sum $\nabla+\theta$ is a well-defined integrable $G$-connection locally.
2.2. Nilpotency
In this subsection, we define nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles. We first make a brief review of distributions and exponentiations. We refer the reader to [Jan87, I.7] for more details about distributions and [Ser94, Sei00] about exponentiations.
Distribution
Let $e\in G$ be the identity element. Denote by $I\subseteq k[G]$ the ideal $I=\{f\in k[G]\text{ }|\text{ }f(e)=0\}$. A distribution on $G$ with support in $e$ of order $\leq n$ is a $k$-linear map $k[G]\rightarrow k$, which maps $I^{n+1}$ to zero. Denote by ${\rm Dist}_{n}(G)$ the set of all distributions on $G$ with support in $e$ of order $\leq n$. Clearly, we have
${\rm Dist}_{n}(G):=(k[G]/I^{n+1})^{*}.$
Let ${\rm Dist}^{+}_{n}(G)\subseteq{\rm Dist}_{n}(G)$ consists of all distributions which send $e$ to zero, and then
${\rm Dist}^{+}_{n}(G)=(I/I^{n+1})^{*},\,{\rm Dist}_{n}(G)\cong k\oplus{\rm Dist}^{+}_{n}(G).$
There is a natural way to define a multiplication on distributions such that
${\rm Dist}_{n}(G)\cdot{\rm Dist}_{m}(G)\subseteq{\rm Dist}_{n+m}(G).$
Then ${\rm Dist}(G):=\bigcup_{n\geq 0}{\rm Dist}_{n}(G)$ is a $k$-algebra, which is called the algebra of distributions on $G$. Moreover, given $f\in{\rm Dist}_{n}(G)$ and $g\in{\rm Dist}_{m}(G)$, we have
$[f,g]=fg-gf\in{\rm Dist}_{n+m-1}(G).$
In particular, if $f,g\in{\rm Dist}^{+}_{1}(G)$, we get
$[f,g]=fg-gf\in{\rm Dist}^{+}_{1}(G),$
which gives a Lie algebra structure on ${\rm Dist}^{+}_{1}(G)$, and ${\rm Dist}^{+}_{1}(G)\cong\mathfrak{g}$. Therefore, given elements $a,b\in\mathfrak{g}$, the multiplication $ab$ is defined as the multiplication of the corresponding distributions. As a special case, the $p$-th power of an arbitrary element $a\in\mathfrak{g}\cong{\rm Dist}^{+}_{1}(G)$ also lies in $\mathfrak{g}$, and we denote it by $a^{[p]}$. If $a^{[p]}=0$, $a$ is called $[p]$-nilpotent. Moreover, if $a$ is $[p]$-nilpotent, then $a$ is nilpotent of exponent $\leq p-1$ because $({\rm ad}(a))^{p}=0$ in characteristic $p$, where ${\rm ad}(a):\mathfrak{g}\rightarrow\mathfrak{g}$ is the adjoint operator.
Exponentiation
Given a connected reductive split reductive linear algebraic group $G$, we suppose that the characteristic $p$ is pretty good. We refer the reader to [Her13, Definition 2.11] for the definition and [Sob15, Introduction] for examples. Roughly speaking, if $G$ is not of type $A$, $p$ is pretty good if $p>7$, while if $G$ is of type $A_{n}$, $p$ should satisfy the conditions that $p>2$ and $p$ is not divided of $n$. Sobaje studied the exponentiation of commuting nilpotent varieties and introduced the following definition:
Definition 2.4 (Definition 2.3 in [Sob15]).
Let $H$ be an affine algebraic group over $k$. We say that $H$ has an exponential map if there exists an $H$-equivariant injective map $\overline{\rm exp}:\mathcal{N}_{1}(\mathfrak{h})\rightarrow H$, where $\mathcal{N}_{1}(\mathfrak{h})$ is the subvariety of $[p]$-nilpotent element in the Lie algebra $\mathfrak{h}$, such that for each $a\in\mathcal{N}_{1}(\mathfrak{h})$, we have
  1. (1)
    the map $\overline{\rm exp}$, which sends $t\in\mathbb{G}_{a}$ to $\overline{\rm exp}(ta)$, is a one-parameter additive subgroup of $H$ whose differential sends $\frac{\partial}{\partial t}$ to $a$;
  2. (2)
    the one-parameter subgroup $\overline{\rm exp}(\mathbb{G}_{a})$ acts trivially (as adjoint action) on $C_{\mathfrak{h}}(a):=\{b\,|\,[a,b]=0\}\subseteq\mathfrak{h}$, i.e.
    ${\rm Ad}(\overline{\rm exp}(a))(b)=b$
    for $b\in C_{\mathfrak{h}}(a)$.
Based on this definition, Sobaje proved the following theorem.
Theorem 2.5 (Theorem 3.4 in [Sob15]).
Let $\mathcal{N}_{1}(\mathfrak{g})$ be the $[p]$-nilpotent variety of $\mathfrak{g}$ and let $\mathcal{U}_{1}(G)$ be the $p$-unipotent variety of $G$. Then there is a unique $G$-equivariant bijection
Indeed, when the nilpotency of $U$ is less than $p$, the $P$-equivariant isomorphism $\mathfrak{u}\xrightarrow{\cong}U$ is first given by Serre [Ser94, §4] and then studied in [Sei00, Proposition 5.2] and [McN02, §8]. The exponential map given by Sobaje is crucial when we studied Cartier and inverse Cartier transform for nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles in this paper.
Nilpotent G-Higgs bundles and nilpotent flat G-bundles
Let $U$ be an open affine subset of $X$ such that $E|_{U}$ is a trivial $G$-bundle. Taking a section $\partial$ of $T_{U}$, $\theta|_{U}(\partial)$ is a section of $E|_{U}(\mathfrak{g})$. Since $E|_{U}(\mathfrak{g})\cong U\times\mathfrak{g}$, $\theta|_{U}(\partial)$ is regarded as an element in $\mathfrak{g}(\mathcal{O}_{U})$. A $G$-Higgs bundle $(E,\theta)$ is nilpotent of exponent $\leq p-1$, if there exists a covering of $X$ by open affine subsets $U$ such that for all local sections $\partial_{1},\dots,\partial_{p}$ of $T_{U}$, we have
$\theta|_{U}(\partial_{1})\cdots\theta|_{U}(\partial_{p})=0,$
where the product is defined as the multiplication of distributions. Let ${\rm HIG}_{p-1}(X/k,G)$ be the category of nilpotent $G$-Higgs bundles on $X$ of exponent $\leq p-1$. Furthermore, if $G={\rm GL}_{n}(k)$, we use the notation ${\rm HIG}_{p-1}(X/k)$ for the category of nilpotent Higgs bundles on $X$ of exponent $\leq p-1$.
Let $(V,\nabla)$ be a flat $G$-bundle on $X$. For any element $v\in T_{X}$, we define a map of sheaves $T_{X}\rightarrow\mathcal{E}nd(\mathcal{O}_{V})$ as $v\rightarrow\nabla(v)^{p}-\nabla(v^{[p]})$, where $v^{[p]}$ is $p$-th power of $v$. By [Kat70, §5], we obtain a morphism $\psi(\nabla):F^{*}_{X/k}T_{X^{\prime}}\rightarrow\mathcal{E}nd(\mathcal{O}_{V})$ of $\mathcal{O}_{X}$-modules, and for any $v\in F^{*}_{X/k}T_{X^{\prime}}$, the following diagram commutes by ($\ast$2.1).
${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$$\scriptstyle{a}$$\scriptstyle{\psi(\nabla)(v)}$$\scriptstyle{\psi(\nabla)(v)\otimes 1+1\otimes\psi(\nabla_{G})(v)}$$\scriptstyle{a\otimes 1}$
This morphism $\psi(\nabla)$ is called the $p$-curvature of $(V,\nabla)$. If there is no ambiguity, we use the notation $\psi$ for the $p$-curvature. Since $\psi(\nabla_{G})=0$, the $p$-curvature $\psi(\nabla)$ can be regarded as a section of $V(\mathfrak{g})\otimes F^{*}_{X/k}\Omega_{X^{\prime}}$, which is called a $F$-Higgs field (see [CZ15, A.6]).
Given a flat $G$-bundle $(V,\nabla)$, it is nilpotent of exponent $\leq p-1$ if there exists a covering of $X$ by open affine subsets $U$ such that for all local sections $\partial_{1},\dots,\partial_{p}$ of $F^{*}_{U^{\prime}/k}T_{U^{\prime}}$, we have
$\psi|_{U}(\partial_{1})\cdots\psi|_{U}(\partial_{p})=0.$
Let ${\rm MIC}_{p-1}(X/k,G)$ be the category of nilpotent flat $G$-bundles on $X$ of exponent $\leq p-1$. Furthermore, if $G={\rm GL}_{n}(k)$, we use the notation ${\rm MIC}_{p-1}(X/k)$ for the category of nilpotent flat $G$-bundles of exponent $\leq p-1$.
Remark 2.6.
The exponent of nilpotency for $G$-Higgs bundles and flat $G$-bundles considered in this paper is only defined for the integer $p$. Now we consider a special case that the exponent is equal to zero. In this case, a nilpotent Higgs field is trivial. Thus, a nilpotent Higgs bundle of exponent $\leq 0$ is exactly a $G$-bundle, and then the category ${\rm HIG}_{0}(X/k,G)$ of nilpotent $G$-Higgs bundles of exponent $\leq 0$ is exactly the category of $G$-bundles. Similarly, a nilpotent flat $G$-bundle of exponent $\leq 0$ implies that it has trivial $p$-curvature. Thus, the category ${\rm MIC}_{0}(X/k,G)$ of nilpotent flat $G$-bundles of exponent $\leq 1$ is the category of flat $G$-bundles with trivial $p$-curvatures.
2.3. Stability Conditions
@@
In this subsection, we suppose that $X$ is a smooth projective variety and we fix an ample line bundle $H$ on $X$. Then the degree of a vector bundle $E$ on $X$ is given with respect to $H$ and denote it by $\deg_{H}E$. If there is no ambiguity, we use the notation $\deg E$ and omit the subscript $H$.
G-bundles
Ramanathan studied stability conditions of principal bundles on curves in characteristic zero [Ram75, Ram96a, Ram96b]. We generalize his definition to smooth projective varieties in characteristic $p$, and we call it the $R$-stability condition. Let $P$ be a parabolic subgroup of $G$ together with a reduction of structure group $\sigma:X\rightarrow E/P$. Then, we get a $P$-bundle $E_{\sigma}$ in the following way
@@
Given a character $\chi:P\rightarrow\mathbb{G}_{m}$, we obtain a line bundle $\chi_{*}E_{\sigma}$ on $X$ and denote it by $L(\sigma,\chi)$.
@@
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center (we refer the reader to [Ram75] for the definition of anti-dominant characters);
  • for any reduction of structure group $\sigma:X\rightarrow E/P$,
@@
G-Higgs bundles
Now we consider the $R$-stability condition of $G$-Higgs bundles and we refer the reader to [KSZ24] for the case of curves. Let $\sigma:X\rightarrow E/P$ be a reduction of structure group. We obtain a $P$-bundle $E_{\sigma}$ together with a map $E_{\sigma}\rightarrow E$. Therefore, we obtain a natural map
$E_{\sigma}(\mathfrak{p})\rightarrow E({\mathfrak{g}}).$
We say that $\sigma$ is $\theta$-compatible if there is a lifting $\theta_{\sigma}:X\rightarrow E_{\sigma}(\mathfrak{p})\otimes\Omega_{X}$ such that the diagram commutes
@@
Recall that a Higgs bundle $(E,\theta)$ is semistable (resp. stable), if for all $\theta$-invariant subbundles $F$, we have
$\frac{\deg F}{{\rm rk}F}\leq\frac{\deg E}{{\rm rk}E}\quad(\text{resp. }<).$. < ) .
Note that we take “$\theta$-invariant subbundles” to define the stability condition of a Higgs bundle. In the case of $G$-Higgs bundles, we choose $\theta$-compatible reductions for the same reason (see [KSZ24, §4]).
Definition 2.9.
@@
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\theta$-compatible reduction of structure group $\sigma:X\rightarrow E/P$,
@@
Denote by ${\rm HIG}^{s(s)}_{p-1}(X/k,G)$ the category of $R$-(semi)stable nilpotent $G$-Higgs bundles of exponent $\leq p-1$, which is a full subcategory of ${\rm HIG}_{p-1}(X/k,G)$.
Flat G-bundles
Let $(V,\nabla)$ be a flat $G$-bundle. Let $\varsigma$ be a reduction of structure group, and then we have a map $V_{\varsigma}\rightarrow V$ (also $\mathcal{O}_{V}\rightarrow\mathcal{O}_{V_{\varsigma}}$). We say that $\varsigma$ is $\nabla$-compatible if the connection $\nabla$ induces a unique connection $\nabla_{\varsigma}:\mathcal{O}_{V_{\varsigma}}\rightarrow\mathcal{O}_{V_{\varsigma}}\otimes\Omega_{X}$ such that the diagram commutes
[…]
When $V$ is a trivial $G$-bundle (or consider $V$ locally), $\nabla$ can be regarded as a section of $V(\mathfrak{g})\otimes\Omega_{X}$ (Example 2.3). In this case, the $\nabla$-compatibility can be defined in the same way as that of $G$-Higgs bundles. More precisely, a $\nabla$-compatible reduction $\varsigma:X\rightarrow V/P$ is equivalent to the condition that there exists a connection $\nabla_{\varsigma}$ of $V_{\varsigma}$ such that the diagram commutes
[…]

### v2

@@ 2.1. G-Higgs bundles and flat G-bundles
where $a$ is the co-action map. Moreover, for each $v\in F^{*}_{X/k}T_{X^{\prime}}$, the $p$-curvature $\psi$ of $\nabla$ induces the following commutative diagram
${\mathcal{O}_{E}}$${\mathcal{O}_{E}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$${\mathcal{O}_{E}}$${\mathcal{O}_{E}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$$\scriptstyle{a}$$\scriptstyle{\psi(v)}$$\scriptstyle{\psi(v)\otimes 1}$$\scriptstyle{a}$
and we refer the reader to [CZ15, A.6] for more details.
@@
In the following, we consider stability conditions and suppose that $X$ is a smooth projective variety. We fix an ample line bundle $H$ on $X$. The degree of a vector bundle $E$ on $X$ is given with respect to $H$ and denote it by $\deg_{H}E$. If there is no ambiguity, we use the notation $\deg E$ and omit the subscript $H$. We briefly review the stability condition for principal bundles on higher dimensional varieties (see [Ram79, Definition 3], [AB01, Definition 1.1] and [GLSS08, §3.2] for instance), which is first considered by Ramanathan [Ram75, Ram96a, Ram96b]. In this paper, this stability condition is called the $R$-stability condition.
Let $P$ be a parabolic subgroup of $G$ together with a reduction of structure group $\sigma:X^{\rm big}\rightarrow(E|_{X^{\rm big}})/P$, where $X^{\rm big}\subseteq X$ is a big open subset. In this paper, a big open subset $X^{\rm big}\subseteq X$ means that ${\rm codim}(X\backslash X^{\rm big})\geq 2$. We get a $P$-bundle $E_{\sigma}$ in the following way
@@
Given a character $\chi:P\rightarrow\mathbb{G}_{m}$, we obtain a line bundle $\chi_{*}((E|_{X^{\rm big}})_{\sigma})$ on $X^{\rm big}$ and denote it by $L(\sigma,\chi)$.
@@
  • for any maximal parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ trivial on the center;
  • for any big open subset $X^{\rm big}\subseteq X$;
  • for any reduction of structure group $\sigma:X^{\rm big}\rightarrow(E|_{X^{\rm big}})/P$,
@@
Now we consider the $R$-stability condition of $\lambda$-connections $(E,\nabla)$. To simplify the notation, we give the following definition for reductions of structure group on $X$ rather than on $X^{\rm big}$. Let $\sigma:X\rightarrow E/P$ be a reduction of structure group. We obtain a $P$-bundle $E_{\sigma}$ and two natural maps
$E_{\sigma}(\mathfrak{p})\rightarrow E({\mathfrak{g}}),\quad{\rm At}(E_{\sigma})\rightarrow{\rm At}(E).$
The reduction of structure group $\sigma$ is said to be $\nabla$-compatible if there is a $\lambda$-connection $\nabla_{\sigma}:T_{X}\rightarrow{\rm At}(E_{\sigma})$ such that the diagram commutes
[…]
Definition 2.5.
A $\lambda$-connection $(E,\nabla)$ is $R$-semistable (resp. $R$-stable), if
  • for any maximal parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ trivial on the center;
  • for any big open subset $X^{\rm big}\subseteq X$;
  • for any $\nabla$-compatible reduction of structure group $\sigma:X^{\rm big}\rightarrow(E|_{X^{\rm big}})/P$,
we have
@@
We give an equivalent definition of the $R$-stability condition in the case of $G$-Higgs bundles (00-connections). Let $(E,\theta)$ be a $G$-Higgs bundle. A reduction of structure group $\sigma$ is said to be $\theta$-compatible if there is a lifting $\theta_{\sigma}:X\rightarrow E_{\sigma}(\mathfrak{p})\otimes\Omega_{X}$ such that the diagram commutes
[…]
@@
  • for any maximal parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ trivial on the center;
  • for any big open subset $X^{\rm big}\subseteq X$;
  • for any $\nabla$-compatible reduction of structure group $\sigma:X^{\rm big}\rightarrow(E|_{X^{\rm big}})/P$,
@@
By Remark 2.3, when $\lambda=0$, a $\nabla$-compatible reduction is equivalent to a $\theta$-compatible reduction. Therefore, the above definition is equivalent to Definition 2.5.
2.2. Nilpotency
We first review the nilpotency for Higgs bundles and flat bundles.
Definition 2.7 ([LSZ15, P.860]).
A Higgs bundle $(E,\theta)$ (resp. flat bundle $(V,\nabla)$) is said to be nilpotent of exponent $\leq n$ if there exists a covering of $X$ by open affine subsets $U_{i}$ such that for each $U_{i}$ and arbitrary sections $\partial_{1},\dots,\partial_{n+1}$ of $T_{U_{i}}$ (resp. $F^{*}_{U^{\prime}_{i}/k}T_{U^{\prime}_{i}}$), we have
$\theta|_{U_{i}}(\partial_{1})\cdots\theta|_{U_{i}}(\partial_{n+1})=0\ (\text{resp.}\ \psi|_{U_{i}}(\partial_{1})\cdots\psi|_{U_{i}}(\partial_{n+1})=0).$
Here is an equivalent description of the nilpotency.
Proposition 2.8.
A rank $r$ Higgs bundle $(E,\theta)$ (resp. flat bundle $(V,\nabla)$) is nilpotent of exponent $\leq n$ if and only if there exists a covering of $X$ by affine open subsets $U_{i}$ such that for each $U_{i}$, the set
$\{\theta|_{U_{i}}(\partial)\ |\ \partial\ \text{is a section of }T_{U_{i}}\}\quad(resp.\text{ }\{\psi|_{U_{i}}(\partial)\ |\ \partial\ \text{is a section of }F^{*}_{U^{\prime}_{i}/k}T_{U^{\prime}_{i}}\})$
lies in $\mathfrak{u}(\mathcal{O}_{U_{i}})$, where $\mathfrak{u}$ is the Lie algebra of a unipotent radical $U$ of a parabolic subgroup $P$ of ${\rm GL}_{r}$ such that the nilpotency class of $U$ is smaller than $n+1$.
The proof of this proposition follows directly from the following result in linear algebra.
Lemma 2.9.
Let $\{A_{1},\dots,A_{m}\}$ be a set of pairwise commutative nilpotent matrices in $\mathfrak{gl}_{r}$. The equality
$A_{1}^{k_{1}}\cdots A_{m}^{k_{m}}=0$
hold for any nonnegative integers $k_{1},\dots,k_{m}$ such that $k_{1}+\dots+k_{m}=n+1$ if and only if $\{A_{1},\dots,A_{m}\}$ lies in the Lie algebra $\mathfrak{u}$ of a nilpotent group $U$ with nilpotency class smaller than $n+1$, where $U$ is the unipotent radical of a parabolic subgroup $P$ of ${\rm GL}_{r}$.
Proof.
We first assume that $\{A_{1},\dots,A_{m}\}$ lies in such a Lie algebra $\mathfrak{u}$. Without loss of generality, we may assume that the parabolic subgroup $P$ is in the standard form, that is, elements of $P$ are of
$\begin{pmatrix}A_{11}&A_{12}&A_{13}&\cdots&A_{1k}\\ 0&A_{22}&A_{23}&\cdots&A_{2k}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&A_{kk}\\ \end{pmatrix},$
where $A_{ij}$ is a $d_{i}\times d_{j}$ block. The unipotent radical $U$ of $P$ is
$\begin{pmatrix}I_{d_{1}}&A_{12}&A_{13}&\cdots&A_{1k}\\ 0&I_{d_{2}}&A_{23}&\cdots&A_{2k}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&I_{d_{k}}\\ \end{pmatrix},$
together with its Lie algebra $\mathfrak{u}$
$\begin{pmatrix}0&A_{12}&A_{13}&\cdots&A_{1k}\\ 0&0&A_{23}&\cdots&A_{2k}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&0\\ \end{pmatrix}.$
Then the equality
$A_{1}^{k_{1}}\cdots A_{m}^{k_{m}}=0$
holds for any nonnegative integers $k_{1},\dots,k_{m}$ such that $k_{1}+\dots+k_{m}\geq k$ since $A_{i}\in\mathfrak{u}$. Since the nilpotency class of $U$ is smaller than $n+1$, we have $k\leq n+1$. Thus, the equality
$A_{1}^{k_{1}}\cdots A_{m}^{k_{m}}=0$
holds for any nonnegative integers $k_{1},\dots,k_{m}$ such that $k_{1}+\dots+k_{m}=n+1$.
For the other direction, let $\{A_{1},\dots,A_{m}\}$ be a set of pairwise commutative nilpotent matrices such that the equalities
$A_{1}^{k_{1}}\cdots A_{m}^{k_{m}}=0$
hold for all nonnegative integers $k_{1},\dots,k_{m}$ such that $k_{1}+\dots+k_{m}=n+1$. We need to find a parabolic subgroup $P$ of ${\rm GL}_{r}$ with unipotent radical $U$, of which the nilpotency class smaller than $n+1$. Indeed, this was proved in [Kos93, P.165-166] and we include the proof here for completeness. Consider each $A_{i}$ as an endmorphism of $V=k^{r}$, and define subspaces
$V_{i}:=\mathop{\bigcap}\limits_{k_{1}+\dots+k_{m}=i}\ker(A_{1}^{k_{1}}\cdots A_{m}^{k_{m}})$
for $10\leq i\leq n+$ 1. Then we obtain a flag of $V$:
$0=V_{0}\subset V_{1}\subset\cdots\subset V_{n+1}=V.$
We can choose a basis of $V$ as
$\{\alpha_{11},\dots,\alpha_{1d_{1}};\ \alpha_{21},\dots,\alpha_{2d_{2}};\ \cdots;\ \alpha_{n+1,1},\dots,\alpha_{n+1,d_{n+1}}\}$
such that
$\{\alpha_{11},\dots,\alpha_{1d_{1}};\ \alpha_{21},\dots,\alpha_{2d_{2}};\ \cdots;\ \alpha_{i,1},\dots,\alpha_{i,d_{i}}\}$
is a basis of $V_{i}$ for each $i$. Under this basis, each matrix $A_{i}$ has the form
$\begin{pmatrix}0&A_{12}&A_{13}&\cdots&A_{1,n+1}\\ 0&0&A_{23}&\cdots&A_{2,n+1}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&0\\ \end{pmatrix},$
Now we define the parabolic subgroup $P$ to be
$\begin{pmatrix}A_{11}&A_{12}&A_{13}&\cdots&A_{1,n+1}\\ 0&A_{22}&A_{23}&\cdots&A_{2,n+1}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&A_{n+1,n+1}\\ \end{pmatrix},$
where $A_{ij}$ is a $d_{i}\times d_{j}$ matrix. Cleary, the unipotent radical $U$ of $P$ is
$\begin{pmatrix}I_{d_{1}}&A_{12}&A_{13}&\cdots&A_{1,n+1}\\ 0&I_{d_{2}}&A_{23}&\cdots&A_{2,n+1}\\ \vdots&\vdots&\vdots&\dots&\vdots\\ 0&0&0&\cdots&I_{d_{n+1}}\\ \end{pmatrix},$
of which the nilpotency class is smaller than $n+1$. ∎
Motivated by Proposition 2.8, we introduce the definition of nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles.